# Adversarial bandit optimization for approximately linear functions

### v4

@@ 3.1 Notation
For any $v\in\mathbb{R}^{d}$, let $v^{\perp}$ be the space orthogonal to $v$. Let $\mathbb{S}^{d}_{1}=\{x\mid\|x\|=1\}$. The vector $e_{i}\in\mathbb{R}^{d}$ is a standard basis vector with a value of $1$ in the $i$-th position and 00 in all other positions. $I$ is an identity matrix with dimensionality implied by context.
@@ 4 Main Results
1:$T$, parameters $\eta\in\mathbb{R},\delta\in(0,1)$, $\nu$-normal barrier $\mathcal{R}$ on $con(\mathcal{K})$
@@ 4.1 SCRiBLe with lifting
The algorithm with parameters $\eta=\frac{\sqrt{2\nu\log T}}{2d\sqrt{T}},\delta=\frac{1}{T^{2}}$ guarantees the following expected regret bound
@@
The algorithm with parameters $\eta=\frac{\sqrt{2\nu\ln T}}{2d\sqrt{T}},\delta=\frac{1}{T^{2}}$ ensures that with probability at least $1-\gamma$
@@ 5.1 $\nu$-normal barrier
If $\mathcal{R}$ be a $\nu$-normal barrier for $\Psi\subseteq R^{d}$, then for any $x\in int(\Psi)$ and any $h\in\Psi$, we have
@@ 7 Experiments
We conduct comparative experiments across a range of $\epsilon$ values: 0, 0.25, 0.5, 0.75, and 1, evaluating Algorithm 1, SCRiBLe [2], and SCRiBLe with lifting and increasing learning rates [17].
@@
We generate the artificial data as follows. The input set is constructed such that each element $x\in\mathbb{R}^{d}$ satisfies $\|x\|_{2}<D$. The loss vectors $\theta_{1},...,\theta_{T}$ are $d$-dimensional vectors randomly generated before the experiment, each with a Euclidean norm no greater than $G$. The $\nu$-normal barrier is defined as $\mathcal{R}(x,b)=400\cdot\left(-\log\left(1-\frac{\|x\|^{2}}{D^{2}}\right)-2\nu\log b\right)$ and $\eta$ is defined as $\frac{20\sqrt{2\log T}}{4d\sqrt{T}}$. In our experiments, we set $D=3$, $G=1$, $d=5$, respectively. The perturbation is defined as $\sigma(y_{t})=\epsilon\sin{((y_{t}^{\top}l)\pi)}$, where $l$ is a random vector in $\mathbb{R}^{d}$ (Note that this is not the worst-case perturbation, which may explain why the results do not increase significantly as $\epsilon$ grows). Each algorithm is run with a fixed time horizon of $T=2000$. For each value of $\epsilon$, we repeat the experiment 10 times for each algorithm and compute the average cumulative loss. The x-axis represents the value of $\epsilon$, and the y-axis shows the average cumulative loss. As shown in the figure above, our method generally outperforms SCRiBLe. Compared to Lee et al.’s method, our approach achieves slightly better performance. This indicates that, when facing oblivious adversaries, not using increasing learning rates (i.e., dynamically adjusting the value of $\eta$) has little to no impact on the results(see Fig. 1).
@@ 8 Conclusion
In this work, we investigate the bandit optimization problem involving a class of non-convex and non-smooth loss functions, where in each trial the loss function is the sum of a linear function and a small but arbitrary perturbation chosen after observing the player’s choice. We introduce a novel approach for analyzing high-probability regret and derive both upper and lower bounds for the problem. For future work, we aim to study the bandit optimization problem for a broader class of non-convex and non-smooth loss functions, such as those formed by the sum of a convex function and a perturbation.
@@ 0.A.1 Proof of Lemma 4
$\|h\|_{x}\leq-\nabla\mathcal{R}(x)^{\top}h\leq\|\nabla\mathcal{R}(x)^{\top}h\rvert.$ (23)
@@
$\|\nabla\mathcal{R}(x)^{\top}h\rvert\\ =\|\nabla\mathcal{R}(x)^{\top}(h-x+x)\rvert\\ \leq\|\nabla\mathcal{R}(x)^{\top}(h-x)\rvert+\|\nabla\mathcal{R}(x)^{\top}x\rvert.$ (24)
By Lemma 1, $\|\nabla\mathcal{R}(x)^{\top}(h-x)\rvert+\|\nabla\mathcal{R}(x)^{\top}x\rvert\leq\nu+\|x^{\top}\nabla^{2}\mathcal{R}(x)x\rvert=2\nu$.
@@ 0.A.3 Proof of Lemma 8
Recall for any $\delta\in(0,1)$, $\mathcal{K}_{\delta}=\{x|\frac{1}{1-\delta}x\in\mathcal{K}\}$ and $\mathcal{K^{\prime}_{\delta}}=\{(x,1):x\in\mathcal{K_{\delta}}\}$. Let $x_{\delta}^{*}=\prod_{\mathcal{K}_{\delta}}x^{*}$, by properties of projections, then
@@
Let $\theta_{t}^{\prime}=(\theta_{t},z)$, where $z$ is the $(d+1)$th coordinate of $d\mathbb{E}_{t}[(\theta_{t},0)^{\top}(x_{t}^{\prime}+\mathbf{A}_{t}\mu_{t})\mathbf{A}_{t}^{-1}\mu_{t}]$. From Lemma 6, we know $d\mathbb{E}_{t}[(\theta_{t},0)^{\top}(x_{t}^{\prime}+\mathbf{A}_{t}\mu_{t})\mathbf{A}_{t}^{-1}\mu_{t}]=\theta_{t}^{\prime}$. Since $g_{t}=df(y_{t})\mathbf{A}_{t}^{-1}\mu_{t}=d\theta_{t}^{\prime\top}(x_{t}^{\prime}+\mathbf{A}_{t}\mu_{t})\mathbf{A}_{t}^{-1}\mu_{t}+d\sigma_{t}(y_{t})\mathbf{A}_{t}^{-1}\mu_{t}$, and let $M_{t}=\mathbb{E}_{t}[d\sigma_{t}(y_{t})\mathbf{A}_{t}^{-1}\mu_{t}]$, then $\theta_{t}^{\prime}=\mathbb{E}_{t}[g_{t}]-M_{t}$ and that we have
@@
By Lemma 1 and Lemma 4, we obtain $\|x_{\delta}^{*^{\prime}}\|_{x_{t}^{\prime}}\leq 2\nu$, $\|x_{t}^{\prime}\|_{x_{t}^{\prime}}=\sqrt{\nu}$

### v5

@@ 3.1 Notation
For any $v\in\mathbb{R}^{d}$, let $v^{\perp}$ be the space orthogonal to $v$. Let $\mathbb{S}^{d}_{1}=\{x\mid\|x\|=1\}$. The vector $e_{i}\in\mathbb{R}^{d}$ is a standard basis vector with a value of 111 in the $i$-th position and 0 in all other positions. $I$ is an identity matrix with dimensionality implied by context.
@@ 4 Main Results
[…]
@@ 4.1 SCRiBLe with lifting
Assumption 1
Let $x^{*}=\arg\min_{x\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(x)$, $x_{\delta}^{*}=\Pi_{\mathcal{K}_{\delta}}(x^{*})$, and $x_{\delta}^{*^{\prime}}=(x_{\delta}^{*},1)$. Assume that for a $\nu$-normal barrier $\mathcal{R}$ on $con(\mathcal{K})$, the following inequality holds: $\nabla\mathcal{R}(x^{\prime}_{t})^{\top}(x_{\delta}^{*^{\prime}}-x^{\prime}_{t})\geq-\nu$.
Assumption 1 is introduced to enable the use of Lemma 4 in proving Lemma 8. Under this assumption, the theoretical upper bound can be established. However, when $\nabla\mathcal{R}(x^{\prime}_{t})^{\top}(x_{\delta}^{*^{\prime}}-x^{\prime}_{t})<-\nu$, our method no longer admits the theoretical upper bound.
@@
Assume Assumptions 1. The algorithm with parameters $\eta=\frac{\sqrt{2\nu\log T}}{2d\sqrt{T}}$, $\delta=\frac{1}{T^{2}}$ guarantees the following expected regret bound
@@
Assume Assumptions 1. The algorithm with parameters $\eta=\frac{\sqrt{2\nu\ln T}}{2d\sqrt{T}}$, $\delta=\frac{1}{T^{2}}$ ensures that with probability at least $1-\gamma$
@@ 5.1 $\nu$-normal barrier
If $\mathcal{R}$ be a $\nu$-normal barrier for $\Psi\subseteq R^{d}$, and $x\in int(\Psi)$, $h\in\Psi$ satisfy $\nabla\mathcal{R}(x)^{\top}(h-x)\geq-\nu$, then
@@ 7 Experiments
We conduct comparative experiments across a range of $\epsilon$ values: 0, 0.25, 0.5, 0.75, and 1, evaluating Algorithm 1, SCRiBLe [2], and SCRiBLe with lifting and increasing learning rates [17]. We generate the artificial data as follows. The input set is constructed such that each element $x\in\mathbb{R}^{d}$ satisfies $\|x\|_{2}<D$. The loss vectors $\theta_{1},...,\theta_{T}$ are $d$-dimensional vectors randomly generated before the experiment, each with a Euclidean norm no greater than $G$. The $\nu$-normal barrier is defined as $\mathcal{R}(x,b)=400\cdot\left(-\log\left(1-\frac{\|x\|^{2}}{D^{2}}\right)-2\nu\log b\right)$ and $\eta$ is defined as $\frac{20\sqrt{2\log T}}{4d\sqrt{T}}$. In our experiments, we set $D=3$, $G=1$, $d=5$, respectively. The perturbation is defined as $\sigma(y_{t})=\epsilon\sin{((y_{t}^{\top}l)\pi)}$, where $l$ is a random vector in $\mathbb{R}^{d}$ (Note that this is not the worst-case perturbation, which may explain why the results do not increase significantly as $\epsilon$ grows). Each algorithm is run with a fixed time horizon of $T=2000$. For each value of $\epsilon$, we repeat the experiment 10 times for each algorithm and compute the average cumulative loss. The x-axis represents the value of $\epsilon$, and the y-axis shows the average cumulative loss. As shown in the figure above, our method generally outperforms SCRiBLe. Compared to Lee et al.’s method, our approach achieves slightly better performance. This indicates that, when facing oblivious adversaries, not using increasing learning rates (i.e., dynamically adjusting the value of $\eta$) has little to no impact on the results(see Fig. 1).
@@ 8 Conclusion
In this work, we study bandit optimization with non-convex, non-smooth losses, where each loss consists of a linear term plus a small adversarial perturbation revealed after the action. We propose a new high-probability regret analysis and establish matching upper and lower bounds. Future work includes extending our results to broader loss classes, such as convex losses with perturbations.
@@ 0.A.1 Proof of Lemma 4
$\|h\|_{x}\leq-\nabla\mathcal{R}(x)^{\top}h\leq\lvert\nabla\mathcal{R}(x)^{\top}h\rvert.$ (23)
@@
$\lvert\nabla\mathcal{R}(x)^{\top}h\rvert\\ =\lvert\nabla\mathcal{R}(x)^{\top}(h-x+x)\rvert\\ \leq\lvert\nabla\mathcal{R}(x)^{\top}(h-x)\rvert+\lvert\nabla\mathcal{R}(x)^{\top}x\rvert.$ (24)
By Lemma 1, $\lvert\nabla\mathcal{R}(x)^{\top}(h-x)\rvert+\lvert\nabla\mathcal{R}(x)^{\top}x\rvert\leq\nu+\lvert x^{\top}\nabla^{2}\mathcal{R}(x)x\rvert=2\nu$.
@@ 0.A.3 Proof of Lemma 8
Recall for any $\delta\in(0,1)$, $\mathcal{K}_{\delta}=\{x|\frac{1}{1-\delta}x\in\mathcal{K}\}$ and $\mathcal{K^{\prime}_{\delta}}=\{(x,1):x\in\mathcal{K_{\delta}}\}$. Let $x_{\delta}^{*}=\Pi_{\mathcal{K}_{\delta}}x^{*}$, by properties of projections, then
@@
[…]
@@
By Assumption 1, Lemma 1 and Lemma 4, we obtain $\|x_{\delta}^{*^{\prime}}\|_{x_{t}^{\prime}}\leq 2\nu$, $\|x_{t}^{\prime}\|_{x_{t}^{\prime}}=\sqrt{\nu}$